# Flow Matching Approach to Reaction Coordinate Evaluation and Optimization

### v1

@@ 5.3 FMRC error
Suppose the velocity fields satisfy conditions in Lemma 5.7 and Lemma 5.9, with velocity fields error $\varepsilon,\ \tilde{\varepsilon}$ and Lipschitz constants $L_{s},\ \tilde{L}_{s}$. Suppose the distributions of $(X_{t},X_{t+\tau}),\ (X_{t},\hat{X}_{t+\tau})$ and $(X_{t},X_{t+\tau}),\ (\hat{X}_{t},X_{t+\tau})$ satisfy conditions in Lemma 5.6 with constants $\beta,\tilde{\beta}$. We denote by $H^{1}_{\rho}$ the standard Sobolev space with norm
@@
(13) $\begin{gathered}\|\mathcal{K}_{\tau}-\hat{\mathcal{K}}_{L}\|_{H}:=\sup_{\|g(x)\|_{H^{1}_{\rho_{0}}}\leq 1,\|f(y)\|_{H^{1}_{\rho_{1}}\leq 1}}\left<g,\mathcal{K}_{\tau}f-\hat{\mathcal{K}}_{L}f\right>_{\rho_{0}}\\ \|\mathcal{T}_{\tau}-\hat{\mathcal{T}}_{D}\|_{H}:=\sup_{\|\tilde{g}(y)\|_{H^{1}_{\rho_{1}}}\leq 1,\|\tilde{f}(x)\|_{H^{1}_{\rho_{0}}\leq 1}}\left<\tilde{g},\mathcal{T}_{\tau}\tilde{f}-\hat{\mathcal{T}}_{L}\tilde{f}\right>_{\rho_{1}}\end{gathered}$
where $f,\tilde{g}\in H_{\rho_{1}}^{1},\ \tilde{f},g\in H_{\rho_{0}}^{1}$. The operator errors are
@@ 6 Numerical examples
$V^{\prime}(x_{1},x_{2})=\cos(7\arctan(x_{1},x_{2}))+10(\sqrt{x_{1}^{2}+x_{2}^{2}}-1)^{2}$
@@ Acknowledgments
The authors would like to express their gratitude to
Mingyuan Zhang and Yajun Wang for their insightful discussions.
@@
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\left<g(x),\mathcal{K}_{\tau}h-\hat{\mathcal{K}}_{L}h\right>$
@@
The detailed proof of the condition $\|f\|_{\dot{H}^{1}(\rho(x,y))}\leq 1$ can be transformed into conditions
@@
$\|f\|_{\dot{H}^{1}(\rho(x,y))}\leq 1$ $\Longleftrightarrow\int_{X\times Y}\|\nabla f(x,y)\|^{2}\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longleftarrow\int_{X\times Y}\|\nabla g(x)h(y)\|^{2}\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longleftrightarrow\int_{X\times Y}\|h(y)\nabla g(x)+g(x)\nabla h(y)\|^{2}\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longleftarrow\int_{X\times Y}\left[\|h(y)\nabla g(x)\|^{2}+\|g(x)\nabla h(y)\|^{2}\right]\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 2$
$\Longleftarrow\int_{X\times Y}\|h(y)\nabla g(x)\|^{2}\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 1,\ \int_{X\times Y}\|g(x)\nabla h(y)\|^{2}\rho_{0}(x)\rho_{1}(y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longleftarrow\int_{Y}h(y)^{2}\rho_{1}(y)\mathrm{d}y\int_{X}\|\nabla g(x)\|^{2}\rho_{0}(x)\mathrm{d}x\leq 1,\ \int_{X}g(x)^{2}\rho_{0}(x)\mathrm{d}x\int_{Y}\|\nabla h(y)\|^{2}\rho_{1}(y)\mathrm{d}y\leq 1$
$\Longleftarrow\int_{Y}h(y)^{2}\rho_{1}(y)\mathrm{d}y\leq 1,\ \int_{X}\|\nabla g(x)\|^{2}\rho_{0}(x)\mathrm{d}x\leq 1,\ \int_{X}g(x)^{2}\rho_{0}(x)\mathrm{d}x\leq 1,$
$\qquad\int_{Y}\|\nabla h(y)\|^{2}\rho_{1}(y)\mathrm{d}y\leq 1$
$\Longleftarrow\sqrt{\int\left(g(x)^{2}+\|\nabla g(x)\|^{2}\right)\rho_{0}(x)\mathrm{d}x}\leq 1,\ \sqrt{\int\left(h(x)^{2}+\|\nabla h(x)\|^{2}\right)\rho_{1}(y)\mathrm{d}x}\leq 1$
Utilize the $W_{2}$ error bound from Lemma 5.7, we can obtain

### v2

# Optimal Low-dimensional Approximation of Transfer Operators via Flow Matching: Computation and Error Analysis

@@ 5.3 FMRC error
Suppose the velocity fields satisfy conditions in Lemma 5.7 and Lemma 5.9, with velocity fields error $\varepsilon,\ \tilde{\varepsilon}$ and Lipschitz constants $L_{s},\ \tilde{L}_{s}$. Suppose the distributions of $(X_{t},X_{t+\tau})$ and its approximation $(X_{t},\hat{X}_{t+\tau})$, as well as $(X_{t},X_{t+\tau})$ and $(\hat{X}_{t},X_{t+\tau})$, both satisfy conditions in Lemma 5.6 with constants $\beta,\tilde{\beta}$. We denote by $H^{1}_{\rho}$ the standard Sobolev space with norm
@@
(13) $\begin{gathered}\|\mathcal{K}_{\tau}-\hat{\mathcal{K}}_{L}\|_{H}:=\sup_{\|g(x)\|_{H^{1}_{\rho_{0}}}\leq 1,\|h(y)\|_{H^{1}_{\rho_{1}}\leq 1}}\left<g(x),\left(\mathcal{K}_{\tau}h-\hat{\mathcal{K}}_{L}h\right)(x)\right>_{\rho_{0}}\\ \|\mathcal{T}_{\tau}-\hat{\mathcal{T}}_{D}\|_{H}:=\sup_{\|\tilde{g}(x)\|_{H^{1}_{\rho_{0}}\leq 1},\|\tilde{h}(y)\|_{H^{1}_{\rho_{1}}}\leq 1}\left<\tilde{h}(y),\left(\mathcal{T}_{\tau}\tilde{g}-\hat{\mathcal{T}}_{L}\tilde{g}\right)(y)\right>_{\rho_{1}}\end{gathered}$
where $h,\tilde{h}\in H_{\rho_{1}}^{1},\ \tilde{g},g\in H_{\rho_{0}}^{1}$. The operator errors are
@@ 6 Numerical examples
$V^{\prime}(x_{1},x_{2})=\cos(7\arctan(x_{2},x_{1}))+10(\sqrt{x_{1}^{2}+x_{2}^{2}}-1)^{2}$
@@ Acknowledgments
The authors would like to express their gratitude to Xin Zhang, Mingyuan Zhang and Yajun Wang for their insightful discussions.
@@
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\int_{X}\int_{Y}g(x)h(y)\left(\rho_{0}(x)p_{\tau}(x,y)-\rho_{0}(x)\hat{p}^{L}_{\tau}(x,y)\right)\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\int_{X}g(x)\rho_{0}(x)\left(\int_{Y}\left(h(y)p_{\tau}(x,y)-h(y)\hat{p}^{L}_{\tau}(x,y)\right)\mathrm{d}y\right)\mathrm{d}x$
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\left<g(x),\mathcal{K}_{\tau}h-\hat{\mathcal{K}}_{L}h\right>_{\rho_{0}}$
@@
For the reduced-order Perron-Frobenius (PF) operator, the proof follows the same approach
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\int_{X}\int_{Y}g(x)h(y)\left(\rho_{0}(x)p_{\tau}(x,y)-\rho_{0}(x)\hat{p}^{L}_{\tau}(x,y)\right)\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\int_{Y}h(y)\rho_{1}(y)\int_{X}\frac{\rho_{0}(x)}{\rho_{1}(y)}\left(p_{\tau}(x,y)-\hat{p}^{L}_{\tau}(x,y)\right)g(x)\mathrm{d}x\mathrm{d}y$
$=\frac{1}{\sqrt{\beta}}\sup_{\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1}\left<h(y),\left(\mathcal{T}_{\tau}g-\hat{\mathcal{T}}_{D}g\right)(y)\right>_{\rho_{1}}$
$=\frac{1}{\sqrt{\beta}}\|\mathcal{T}_{\tau}-\hat{\mathcal{T}}_{D}\|_{H}$
The detailed proof that the condition $\|f\|_{\dot{H}^{1}(\rho(x,y))}\leq 1$ can be transformed into the conditions
@@
$\|f\|_{\dot{H}^{1}(\rho(x,y))}\leq 1\Longleftrightarrow\int_{X\times Y}\|\nabla f(x,y)\|^{2}_{\mathbb{R}^{2n}}\rho(x,y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longrightarrow\int_{X\times Y}\|\nabla\cdot\left(g(x)h(y)\right)\|^{2}_{\mathbb{R}^{2n}}\rho(x,y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longrightarrow\int_{X\times Y}\left(\|h(y)\nabla_{x}g(x)\|^{2}_{\mathbb{R}^{n}}+\|g(x)\nabla_{y}h(y)\|^{2}_{\mathbb{R}^{n}}\right)\rho(x,y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longrightarrow\int_{X\times Y}h^{2}(y)\|\nabla_{x}g(x)\|^{2}_{\mathbb{R}^{n}}\rho(x,y)\mathrm{d}x\mathrm{d}y+\int_{X\times Y}g^{2}(x)\|\nabla_{y}h(y)\|^{2}_{\mathbb{R}^{n}}\rho(x,y)\mathrm{d}x\mathrm{d}y\leq 1$
$\Longrightarrow\|g(x)\|_{H^{1}(\rho_{0}(x))}\leq 1,\quad\|h(y)\|_{H^{1}(\rho_{1}(y))}\leq 1$